can you please solve this problems for me
6x=4
3+x=-2
2x+3x=6-2x
2x-4+-4x+2=12
3 5
5x-6-3x+4=7
2 5

Answer by edjones(8007)

6x=4
x=4/6=2/3
.
Ed